# A Bayesian Dirichlet Auto-Regressive Conditional Heteroskedasticity Model for Forecasting Currency Shares

### v2

@@ 1 Introduction
Each day’s currency proportions must lie in a simplex (they are nonnegative and sum to one), a constraint that complicates standard time-series methods. One classical approach is to use log-ratio transformations, as pioneered by Aitchison (1982) and extended in works such as Cargnoni et al. (1997), Ravishanker et al. (2001), and Mills (2010), among others. Transformed additive log-ratio VARMA (tVARMA) models have been fruitfully applied to compositional data (Brunsdon and Smith, 1998; Mills, 2009, 2010).
@@
In contrast, the Dirichlet distribution models compositional data directly on the simplex. Benjamin et al. (2003) proposed frequentist Generalized ARMA models that could be extended to the simplex. Contributions to Dirichlet time series include Bayesian compositional state space models by Grunwald et al. (1993), da Silva et al. (2011), and da Silva and Rodrigues (2015); the Bayesian Dirichlet Auto-Regressive Moving Average (B-DARMA) model by Katz et al. (2024, 2025); Katz and Maierhofer (2025); Katz (2025); and frequentist Dirichlet ARMA models by Zheng and Chen (2017).
Seminal works in Bayesian time series analysis include West (1996), Prado and West (2010), Barber et al. (2011), Berliner (1996), Pole et al. (2018), and Koop and Korobilis (2010). Bayesian methods have been applied to Vector Auto-Regression (VAR) and Vector Auto-Regression Moving Average (VARMA) models (Spencer, 1993; Uhlig, 1997; Bańbura et al., 2010; Karlsson, 2013). Bayesian generalized linear models (GLMs) for time series data, suitable for counts and categorical responses, have been explored by Brandt and Sandler (2012), Roberts and Penny (2002), Nariswari and Pudjihastuti (2019), Chen and Lee (2016), McCabe and Martin (2005), Berry and West (2020), Fukumoto et al. (2019), Silveira de Andrade et al. (2015), and West (2013).
The generalized auto-regressive conditional heteroskedasticity (GARCH) model (Bollerslev, 1986) is fundamental in modeling financial time series volatility. Extensions include non-normal likelihoods to capture heavy tails and skewness (Nelson, 1991; Engle, 2001). Bauwens et al. (2006) surveyed advancements in multivariate GARCH models.
@@
The remainder of the paper is organized as follows. The next section introduces the B-DARCH model. Section 3 reports simulation studies comparing B-DARCH with three alternatives: (i) a B-DARMA data model; (ii) a Bayesian transformed-data normal VARMA model; and (iii) a time-varying-parameter comparator in ALR space (B-TVP-tVARMA). Section 4 analyzes the Airbnb data. The paper concludes with a brief discussion.
@@ 3.5 Fitted Models and Priors
For B‑DARCH, the initial log‑precision $\phi_{0}\sim\mathcal{N}(7,\,1.5^{2})$ and the AR/MA terms in $\log\phi_{t}$ have $\mathcal{N}(0.35,\,0.5^{2})$ and $\mathcal{N}(-0.75,\,0.5^{2})$ priors, respectively.
@@ 4.1.1 Regional Currency Dynamics
The currency proportions for each region are shown in Figures 1 and Supplementary Figures S3 and S4. Using the U.S. Dollar (USD) as the reference category, 30-day rolling ALR means are shown in Supplementary Figures S5 and S6 while 30-day rolling ALR variances are shown in Supplementary Figure S7 and Figure 2. The correlations of the ALR values with their lag 1 values is shown in Supplementary Figure S8.
Across all four regions, currencies exhibit seasonal cycles and trends relative to USD. Regions 1 and 3 display pronounced yearly seasonality for multiple currencies, with Region 4 exhibiting the strongest seasonal patterns. Pre-COVID-19 (before 2020), currencies in Regions 1 and 2 generally showed steady levels or consistent declines relative to USD, while in Region 3, the Australian Dollar (AUD) and other currencies displayed distinct upward trends. During the COVID-19 period, starting around January or February of 2020, significant shifts are seen: in Region 1, all currencies except the Canadian Dollar (CAD) saw large declines in their means at the expense of the USD; in Region 2, the Euro (EUR) and Swiss Franc (CHF) experienced the largest increases as the USD decline; in Region 3, AUD continued its sharp rise, while the New Zealand Dollar (NZD) initially declined but rapidly recovered; in Region 4, the Brazilian Real (BRL) experienced sharp increases in its rolling mean before being overtaken by USD, and the Mexican Peso (MXN) peaked before a drastic decline.
Lag 1 correlations show strong autocorrelations within currencies, with high self-correlations observed in EUR and GBP in Region 1 and in BRL, Chilean Peso (CLP), and MXN in Region 4. Positive lag 1 cross correlations were found among currencies in Region 2, except for CAD, while in Region 3, NZD and AUD were strongly correlated, and EUR was negatively correlated with other currencies. In Region 4, negative correlations between BRL and EUR, as well as between CLP and EUR, suggest opposing movements between these currency pairs.
Figure 2 and Supplementary Figure S7 show the 30-day rolling variance of each currency’s ALR-transformed proportions. Rather than a single abrupt change, we observe recurrent episodes of elevated variability, often persisting for several weeks before subsiding. For instance, Region 1 (top left) shows multiple surges for EUR and GBP, culminating in early 2020 and tapering mid-year, while Region 2 (top right) begins with a high-variance phase (notably in CAD and CHF), settles to moderate levels around 0.1–0.2, and then rises again at the onset of 2020. Region 3 (bottom left) experiences shorter bursts around 2018 and late 2019, each followed by another peak in early 2020, suggesting that once volatility increases, it may linger rather than immediately revert. Finally, Region 4 (bottom right) transitions from moderate day-to-day swings to abrupt climbs in BRL and CLP near 2019, with additional spikes in BRL and MXN emerging after 2020. In every region, these heightened-variance intervals emerge at distinct times and exhibit varying durations, indicating that volatility in currency shares is shaped by more than one-off shocks, and can arise repeatedly under evolving market conditions. Such patterns underscore the importance of accommodating dynamic, recurrent shifts in variance for effective modeling and forecasting.
@@ 4.7.1 Model Specification
Out-of-sample validation results (Supplementary Table S1) show that Region 1 performed best with $K_{\text{year}}=8$ and $(P,Q)=(1,0)$, yielding a average FMAE of 0.0036 and total FRSS of .0173. Region 2 achieved optimal performance with $K_{\text{year}}=10$ and $(P,Q)=(1,1)$, resulting in a average FMAE of .0127 and tRSS of .2018. In Region 3, the optimal model had $K_{\text{year}}=10$ and $(P,Q)=(1,1)$, with a average FMAE of .0283 and total FRSS of 1.1395. Region 4 was best modeled with $K_{\text{year}}=8$ and $(P,Q)=(3,2)$, with a FMAE of .0116 and FRSS of .1553.
@@
Table 4 reports Forecast Mean Absolute Error (FMAE) and Forecast Residual Sum of Squares (FRSS) for each region and currency across four models: B-DARCH, B-DARMA, B-tVARMA, and a time-varying‑parameter VARMA fit in ALR space (TVP-tVARMA). The corresponding out-of-sample forecasts $\bar{\mu}_{tj}$ are shown in Figures 36.
When errors are aggregated over currencies within a region (the total rows of Table 4), B-DARCH attains the lowest total FMAE and FRSS in Regions 1–4. For Region 1, totals are FMAE/FRSS == 2.84/3.15 (B-DARCH) vs. 7.65/19.11 (B-DARMA), 5.20/10.62 (B-tVARMA), and 4.55/8.12 (B-TVP-tVARMA). For Region 2, totals are 13.55/73.93 (B-DARCH) vs. 14.71/81.87 (B-DARMA), 17.13/104.43 (B-tVARMA), and 14.68/90.49 (B-TVP-tVARMA). For Region 3 the ordering is 16.23/110.83 (B-DARCH) << 20.17/159.66 (B-DARMA) << 20.73/143.39 (B-TVP-tVARMA) << 23.70/187.09 (B-tVARMA). For Region 4, B-DARCH again leads with 11.21/32.55, followed by B-TVP-tVARMA (12.22/39.85), B-DARMA (15.25/62.33), and B-tVARMA (16.53/77.64). These results reinforce that explicitly modeling time‑varying volatility via DARCH yields the most accurate aggregate forecasts across regions. Across regions the B-TVP-tVARMA comparator required the longest runtime, followed by B-DARCH, while B-DARMA and B-tVARMA typically completed near one hour (Table S4).
@@
Currency-level minima vary by region. In Region 2, B-DARCH delivers the lowest FMAE for five of six currencies (CAD, CHF, EUR, USD, and other), with B‑TVP‑tVARMA best for GBP. In Region 3, B-DARCH is best for AUD, EUR, and other; B‑TVP‑tVARMA is best for JPY and NZD; and B-DARMA has the smallest USD FMAE. Region 4 shows a similar mix: B-DARCH leads for BRL, CLP, and other; B‑TVP‑tVARMA leads for MXN and USD; B-DARMA leads for EUR. In Region 1, B‑DARCH is best for CAD, USD, and other on both FMAE and FRSS; TVP‑tVARMA is best for EUR and JPY; and GBP is effectively a tie after rounding, with TVP‑tVARMA narrowly ahead on both metrics. Overall, the table indicates that the DARCH volatility component is beneficial across all regions, with especially large aggregate gains where volatility shifts are most pronounced.
Figure 3: Airbnb data analysis — Region 1: 92-day forecasts (blue) with 95% credible intervals (shaded) for each currency from Oct. 1 to Dec. 31, 2020, compared to actual values (red) from the preceding six months. Models shown: B-DARCH, B-DARMA, B-tVARMA, and B-TVP-tVARMA.
Figure 4: Airbnb data analysis — Region 2: 92-day forecasts (blue) with 95% credible intervals (shaded) for each currency from Oct. 1 to Dec. 31, 2020, compared to actual values (red) from the preceding six months. Models shown: B-DARCH, B-DARMA, B-tVARMA, and B-TVP-tVARMA.
Figure 5: Airbnb data analysis — Region 3: 92-day forecasts (blue) with 95% credible intervals (shaded) for each currency from Oct. 1 to Dec. 31, 2020, compared to actual values (red) from the preceding six months. Models shown: B-DARCH, B-DARMA, B-tVARMA, and B-TVP-tVARMA.
Figure 6: Airbnb data analysis — Region 4: 92-day forecasts (blue) with 95% credible intervals (shaded) for each currency from Oct. 1 to Dec. 31, 2020, compared to actual values (red) from the preceding six months. Models shown: B-DARCH, B-DARMA, B-tVARMA, and B-TVP-tVARMA.
@@ Acknowledgement
The authors thank Sean Wilson, Jess Needleman, Liz Medina, and Yuling Kuo for helpful discussions, Adam Liss and Ellie Mertz for championing the research.

### v3

@@ 1 Introduction
Each day’s currency proportions must lie in a simplex (they are nonnegative and sum to one), a constraint that complicates standard time-series methods. One classical approach is to use log-ratio transformations, as pioneered by aitchison1982statistical and extended in works such as Cargnoni1997BayesianFO, ravishanker2001compositional, mills2010forecasting, and snyder2017forecasting, among others. Transformed additive log-ratio VARMA (tVARMA) models have been fruitfully applied to compositional data (brunsdon1998time; mills2009forecasting; mills2010forecasting).
@@
In contrast, the Dirichlet distribution models compositional data directly on the simplex. benjamin2003generalized proposed frequentist Generalized ARMA models that could be extended to the simplex. Contributions to Dirichlet time series include Bayesian compositional state space models by grunwald1993time, da2011dynamic, and da2015bayesian; the Bayesian Dirichlet Auto-Regressive Moving Average (B-DARMA) model by KATZ20241556; forecast7030032; forecast7040062; katz2025centeredmadirichletarma; and frequentist Dirichlet ARMA models by zheng2017dirichlet; creus2021dirichlet.
The generalized auto-regressive conditional heteroskedasticity (GARCH) model (bollerslev1986generalized) is fundamental in modeling financial time series volatility. Extensions include non-normal likelihoods to capture heavy tails and skewness (nelson1991conditional; engle2001theoretical). bauwens2006multivariate surveyed advancements in multivariate GARCH models.
@@
The remainder of the paper is organized as follows. The next section introduces the B-DARCH model. Section 3 reports simulation studies comparing B-DARCH with three alternatives: (i) a B-DARMA data model; (ii) a Bayesian transformed-data normal VARMA model; and (iii) a time-varying-parameter comparator in ALR space (B-TVP-tVARMA). Section 4 analyzes the Airbnb data. The paper concludes with a brief discussion. Additional tables and figures appear in the online Supplementary Material (Section 6).
@@ 3.5 Fitted Models and Priors
For B‑DARCH, the initial log‑precision $\phi_{0}\sim\mathcal{N}(7,\,1.5^{2})$ and the AR/MA terms in $\log\phi_{t}$ have $\mathcal{N}(0.5,\,0.15^{2})$ and $\mathcal{N}(-0.75,\,0.15^{2})$ priors, respectively.
@@ 4.1.1 Regional Currency Dynamics
Three features of the data motivate our modeling choices: pronounced seasonality, strong temporal dependence, and recurrent volatility shifts.
Seasonality and trends.
Figure 1 displays currency proportions across the four regions. All regions exhibit clear yearly seasonal cycles, with Region 4 showing the strongest patterns. Pre-pandemic trends vary by region: Regions 1 and 2 show steady or declining currency shares relative to USD, while Region 3 exhibits upward drift in AUD. The COVID-19 period (early 2020) induced sharp, region-specific disruptions. USD surged in Region 1, EUR and CHF rose in Region 2, and BRL experienced large swings in Region 4. These trend breaks motivate our inclusion of flexible seasonal components and suggest the need for models that are robust to structural shifts.
Temporal dependence.
Lag-1 autocorrelations of ALR-transformed series are uniformly high across regions (Supplementary Figure S8), confirming that currency compositions are persistent and justifying autoregressive structure in our specifications.
Time-varying volatility.
Figure 2 reveals that volatility in ALR-transformed proportions is not constant but exhibits recurrent episodes of elevation. Rather than a single shock, we observe multiple surges across all four regions, often persisting for weeks before subsiding. Region 2 begins with elevated variance in CAD and CHF, settles to moderate levels, then spikes again in early 2020. Region 4 shows abrupt climbs in BRL and CLP near 2019 with additional spikes emerging after 2020. These patterns indicate that volatility is shaped by evolving market conditions rather than one-off events. This observation is the central motivation for our DARCH specification.
@@ 4.7.1 Model Specification
We selected region-specific lag orders and Fourier harmonics via out-of-sample validation, with full results reported in Supplementary Table S1. Optimal specifications varied across regions: Regions 2 and 3 favored richer seasonal structure ($K_{\text{year}}=10$) with ARMA(1,1) dynamics, Region 1 performed best with fewer harmonics ($K_{\text{year}}=8$) and a pure autoregressive specification, while Region 4 required the most complex short-run dynamics at ARMA(3,2) despite fewer harmonics. These differences likely reflect heterogeneity in seasonal persistence and short-run adjustment across geographic markets.
@@
Table 4 compares forecast accuracy across four models: B-DARCH, B-DARMA, B-tVARMA, and a time-varying-parameter variant (TVP-tVARMA). Corresponding forecast trajectories for each region’s dominant currency appear in Figures 36.
The central finding is that B-DARCH achieves the lowest aggregate FMAE and FRSS in all four regions. The gains are most pronounced in Region 1, where B-DARCH reduces total FMAE by over 60% relative to B-DARMA and FRSS by more than 80%. Regions 2–4 show more modest but consistent improvements. These results suggest that explicitly modeling time-varying volatility—the distinguishing feature of the DARCH specification—yields meaningful forecast gains, particularly in markets where volatility shifts are most pronounced.
Among the comparators, TVP-tVARMA generally ranks second, indicating that time-varying parameters provide some benefit even without the DARCH volatility structure. However, this flexibility comes at computational cost: TVP-tVARMA required the longest runtimes, while B-DARMA and B-tVARMA typically completed in approximately one hour (Table S4).
@@
While B-DARCH dominates at the aggregate level, currency-level rankings are more nuanced (Table 4). B-DARCH achieves the lowest FMAE for the majority of individual currencies in each region—for instance, five of six currencies in Region 2. However, TVP-tVARMA occasionally outperforms for specific series, including GBP in Region 2, JPY and NZD in Region 3, and MXN and USD in Region 4. In Region 1, B-DARCH leads for CAD, USD, and the residual category, while TVP-tVARMA edges ahead for EUR, JPY, and GBP. B-DARMA is occasionally competitive at the currency level (e.g., USD in Region 3, EUR in Region 4) but never leads in aggregate.
These patterns suggest that DARCH’s volatility modeling provides broad benefits across currency compositions, while time-varying mean parameters may better capture dynamics for particular series. Readers interested in specific currencies can consult the detailed breakdown in Table 4.
Figure 3: Airbnb data analysis — Region 1: 92-day forecasts (blue) with 95% credible intervals (shaded) for USD currency from Oct. 1 to Dec. 31, 2020, compared to actual values (red) from the preceding six months. Models shown: B-DARCH, B-DARMA, B-tVARMA, and B-TVP-tVARMA.
Figure 4: Airbnb data analysis — Region 2: 92-day forecasts (blue) with 95% credible intervals (shaded) for EUR currency from Oct. 1 to Dec. 31, 2020, compared to actual values (red) from the preceding six months. Models shown: B-DARCH, B-DARMA, B-tVARMA, and B-TVP-tVARMA.
Figure 5: Airbnb data analysis — Region 3: 92-day forecasts (blue) with 95% credible intervals (shaded) for AUD currency from Oct. 1 to Dec. 31, 2020, compared to actual values (red) from the preceding six months. Models shown: B-DARCH, B-DARMA, B-tVARMA, and B-TVP-tVARMA.
Figure 6: Airbnb data analysis — Region 4: 92-day forecasts (blue) with 95% credible intervals (shaded) for BRL currency from Oct. 1 to Dec. 31, 2020, compared to actual values (red) from the preceding six months. Models shown: B-DARCH, B-DARMA, B-tVARMA, and B-TVP-tVARMA.
@@ Acknowledgement
The authors thank Sean Wilson, Jess Needleman, Liz Medina, Yuling Kuo, Erica Savage, and Jackson Wang for helpful discussions, Adam Liss and Ellie Mertz for championing the research.
@@
6 Supplementary Materials
Alternative Log-Ratio Transformations
Although we have used the additive log-ratio (ALR) link in equations (4), (5), and (6), other transformations to and from the simplex may be employed. A common choice is the centered log-ratio (clr), defined as
${\textnormal{clr}}(\mathbf{y})\;=\;\Bigl(\,\ln\bigl(\tfrac{y_{1}}{g(\mathbf{y})}\bigr),\;\ln\bigl(\tfrac{y_{2}}{g(\mathbf{y})}\bigr),\;\dots,\;\ln\bigl(\tfrac{y_{J}}{g(\mathbf{y})}\bigr)\Bigr),$
where $g(\mathbf{y})=(\,y_{1}\,y_{2}\,\cdots\,y_{J})^{1/J}$ is the geometric mean of $\mathbf{y}$. The clr mapping places all $J$ components in a transformed $\mathbb{R}^{J}$ space but introduces the constraint that their sum is zero.
Another option is to use the Isometric Log-Ratio (ilr) transformation. For each index $j$, the ilr-transformed coordinate can be expressed as
$z_{j}=\sqrt{\frac{r_{j}}{r_{j}+1}}\,\ln\biggl(\frac{g_{j}(\mathbf{y)}}{\bigl(\prod_{i\in H_{j}}y_{i}\bigr)^{1/r_{j}}}\biggr),$
where $g_{j}(\mathbf{y})$ is defined as the geometric mean of a chosen subset $S_{j}\subseteq\mathbf{y}$, and $H_{j}$ is the complement of $S_{j}$ in $\mathbf{y}$. The value $r_{j}$ is the number of elements in $H_{j}$. The particular way in which $S_{j}$ and $H_{j}$ are selected depends on the specific application and how one wishes to interpret the ilr coordinates.
These mappings are mutually related by linear transformations (egozcue2003isometric) and each provides a bijection from the (J1)(J-1)-dimensional simplex onto a (J1)(J-1)-dimensional subset of $\mathbb{R}^{J}$ or $\mathbb{R}^{J-1}$. Therefore, if the same priors and model assumptions are consistently adapted to each log-ratio space, the resulting compositional models remain equivalent—differing only in how one embeds the data in a Euclidean domain.
Definition of the Partial Autocorrelation Function (PACF)
The PACF at lag $k$ measures the partial correlation between the SSR at time $t$ and the SSR at time tkt-k, controlling for the effects of intermediate lags $1$ to k1k-1. It is defined as the last coefficient $\phi_{kk}$ in the autoregressive model of order $k$
$\text{SSR}_{t}=\phi_{k1}\text{SSR}_{t-1}+\phi_{k2}\text{SSR}_{t-2}+\dots+\phi_{k,k-1}\text{SSR}_{t-(k-1)}+\phi_{kk}\text{SSR}_{t-k}+\varepsilon_{t}$ (8)
where $\varepsilon_{t}$ is the white noise error term at time $t$. The coefficient $\phi_{kk}$ represents the PACF at lag $k$ and can be calculated recursively using the Yule-Walker equations
$\phi_{kk}=\frac{r_{k}-\sum_{j=1}^{k-1}\phi_{k-1,j}r_{k-j}}{1-\sum_{j=1}^{k-1}\phi_{k-1,j}r_{j}}$
Here, $r_{k}$ is the autocorrelation at lag $k$, and $\phi_{k-1,j}$ are the PACF coefficients from the previous lag k1k-1.
By examining the PACF of the SSRs, we can identify the presence of residual autocorrelation at various lags, which indicates whether the model has adequately captured the temporal dependencies in the data.
Standardized Residuals
For the B-DARMA and B-DARCH models, which are based on the Dirichlet distribution as defined in Section 2, we compute standardized residuals using the standard deviations derived from the Dirichlet parameters.
At each time t, the residuals are calculated as $\text{residual}_{tj}=y_{tj}-\mu_{tj},\quad\text{for }j=1,\dots,J.$ The variance of each component $y_{tj}$ under the Dirichlet distribution, Dirichlet($\phi_{t}\boldsymbol{\mu}_{t}$), is $\operatorname{Var}[y_{tj}]=\frac{\mu_{tj}(1-\mu_{tj})}{\phi_{t}+1}$, with standard deviation $\sigma_{tj}=\sqrt{\frac{\mu_{tj}(1-\mu_{tj})}{\phi_{t}+1}}.$
The standardized residuals are then computed as $\text{standardized\_residual}_{tj}=\frac{y_{tj}-\mu_{tj}}{\sigma_{tj}}$ and the sum of squared standardized residuals at time t is $\text{SSR}_{t}=\sum_{j=1}^{J}\left(\text{standardized\_residual}_{tj}\right)^{2}.$
For the B-tVARMA and B-TVP-tVARMA models, residuals are computed in the additive log-ratio (alr) transformed space, $\text{residual}_{tj}={\textnormal{alr}}(\mathbf{y}_{tj})-{\textnormal{alr}}(\mathbf{\hat{y}}_{tj}),$ where $j=1,\dots,4$. These residuals are standardized using the inverse of the Cholesky factor $L_{\Sigma}$ of the estimated covariance matrix $\Sigma$$\text{standardized\_residual}_{t}=L\Sigma^{-1}\times\text{residual}_{t}$. The sum of squared standardized residuals at time t is then calculated as $\text{SSR}_{t}=\sum_{j=1}^{J-1}(\text{standardized\_residual}_{tj})^{2}.$
By examining the PACF of $\{\text{SSR}\}_{t=1}^{T}$ we assess the presence of autocorrelation in the residuals. A well-specified model should produce residuals with minimal autocorrelation, indicating that temporal dependencies have been adequately captured.
Parameter Estimates
Supplementary Figure S13 shows the posterior densities of the auto-regressive ($\alpha$) and moving average ($\tau$) coefficients for each region. In Region 1, the posterior of $\alpha$ is concentrated between 0.25 and 0.75, peaking around 0.45, indicating moderate volatility persistence; shocks have a noticeable but not prolonged effect on volatility. In Region 2, the $\alpha$ coefficient has posterior mean around 0.8, reflecting strong dependence of current volatility on past volatility, and the $\tau$ coefficient centers around 2.75-2.75, indicating that large deviations from expected currency proportions significantly increase / decrease future volatility—characteristic of volatility clustering. Region 3 shows $\alpha$ tightly concentrated between 0.85 and 0.90 and $\tau$ centered around 1.2-1.2 suggesting very strong volatility persistence. In contrast, Region 4 exhibits complex dynamics: $\alpha_{1}$ is centered around $0.15$, indicating small persistence from the first lag; $\alpha_{2}$ is skewed positive and centered around $0.35$, suggesting a stronger influence from the second lag; $\alpha_{3}$ peaks slightly above 0.5, indicating moderate persistence; and the $\tau$ parameter is concentrated near $0.02$, implying minimal influence of past innovations.
Overall, the higher $\alpha$ values in Regions 2 and 3 indicate that volatility in currency fee proportions is more persistent in these regions. The negative $\tau$ coefficients suggest that large deviations—such as those caused by abrupt travel restrictions, policy changes, and fluctuating foreign exchange (FX) rates lead to increased future volatility. This implies that Regions 2 and 3 are more susceptible to external shocks affecting travel and currency markets, resulting in prolonged volatility in currency fee proportions. In contrast, Region 1, with moderate volatility persistence, tends to return to stability more rapidly after shocks, possibly due a more stable currency environment or a larger share of domestic travel. Region 4’s dynamics, characterized by less emphasis on immediate past volatility and innovations, may be attributed to regional factors such as diversified travel sources, differing policy responses, or economic structures.
Figure S1: Average PACF of the sum of squared standardized residual values across all 50 simulations for studies 1 & 2. B-DARCH consistently exhibits lower PACF values beyond lag 1, indicating reduced residual autocorrelation and suggesting that its dynamic precision component better captures volatility shocks in these simulated compositional time series.
Figure S2: Average PACF of the sum of squared standardized residual values across all 50 simulations for studies 3 & 4. The B-DARCH model (red) again shows smaller autocorrelation at higher lags, whereas B-DARMA (green) and B-tVARMA (red) leave more persistent structures in the residuals. This result underlines the advantage of modeling time-varying volatility when faced with abrupt changes in compositional data.
Figure S3: Airbnb data analysis- Proportion of fees by currency for four regions: weekly seasonal behavior. AUD is the Australian dollar, BRL is the Brazillian Real, CAD is the Canadian Dollar,CHF is the Swiss Franc, CLP is the Chilean Peso,EUR is the European Euro, GBP is the Great British Pound, MXN is the Mexican Peso, NZD is the New Zealand Dollar, and USD is the US Dollar. Strong weekly cycles appear in many currencies, motivating the inclusion of weekly seasonal terms in our model’s mean structure.
Figure S4: Airbnb data analysis - proportion of fees by billing currency for four regions from Jan 1, 2020 to June 30, 2020. AUD is the Australian dollar, BRL is the Brazillian Real, CAD is the Canadian Dollar,CHF is the Swiss Franc, CLP is the Chilean Peso,EUR is the European Euro, GBP is the Great British Pound, MXN is the Mexican Peso, NZD is the New Zealand Dollar, and USD is the US Dollar. Distinct spikes and dips across multiple currencies (e.g., sharp increase in BRL or EUR) reflect significant disruptions.
Figure S5: Airbnb data analysis - proportion of fees by currency- 30-day rolling ALR means for four regions from Jan 1, 2020 to June 30, 2020. USD is the reference component.
Figure S6: Airbnb data analysis - proportion of fees by currency- 30-day rolling ALR means for four regions from Jan 1, 2017 to June 30, 2020. USD is the reference component. Apart from seasonal patterns, we see notable multi-year trends and cyclical behaviors (e.g., increasing share of EUR in Region 2, upward drift in AUD in Region 3), underlining the presence of both trend and seasonality in compositional dynamics.
Figure S7: Airbnb data analysis - proportion of fees by currency- 30-day rolling ALR variance for four regions from Jan 1, 2017 to June 30, 2020. USD is the reference component.The data display periods of high variance that sometimes persist for weeks, demonstrating “volatility clustering.” These episodes justify the GARCH-like specification in our B-DARCH model, which allows the precision parameter $\phi_{t}$ to track spikes in uncertainty over time.
Figure S8: Airbnb data analysis- correlation between ALR and lagged ALR values for each currency in the four regions. USD is the reference component. High positive autocorrelations in many currencies (e.g., EUR, GBP) highlight the strong temporal persistence. Negative cross-lag correlations for some pairs (e.g., EUR vs. BRL in Region 4) indicate that certain currencies move in opposite directions. Both effects reinforce the need for a multi-component time-series framework with auto- and cross-dependence.
Figure S9: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 1.
Figure S10: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 2.
Figure S11: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 3.
Figure S12: Airbnb data analysis: Partial Autocorrelation Function (PACF) values for the sum of squared standardized residuals (SSR) on the test set from October 1, 2020, to December 31, 2020, for Region 4.
Figure S13: Airbnb data analysis: Densities of DARCH model coefficients for all 4 regions. Region 1 was modeled with a DARCH(1,0), Regions 2-3 modeled with a DARCH(1,1), and Region 4 modeled with a DARCH(3,1). Each plot shows the estimated density of a specific $\alpha_{l}$ (ar_phi) or $\tau_{k}$ (ma_phi) parameter. Regions 2 and 3 exhibit high $\alpha$ indicating persistent volatility once spikes occur, whereas Region 4’s coefficients suggest moderate or delayed response